As Reported Annual Balance Sheet

Report Date

10/31/2018

Currency

USD

Audit Status

Not Available

Consolidated

Yes

Scale

Millions

Cash & cash equivalents

5166

Accounts receivable, gross

5242

Allowance for doubtful accounts

129

Accounts receivable

5113

Finished goods

4019

Purchased parts & fabricated assemblies

2043

Inventory

6062

Value-added taxes receivable

865

Available-for-sale investments

711

Supplier & other receivables

2025

Prepaid & other current assets

1445

Other current assets

5046

Total current assets

21387

Land, buildings & leasehold improvements

1893

Machinery & equipment, including equipment held for lease

4216

Property, plant & equipment, gross

6109

Accumulated depreciation

3911

Property, plant & equipment

2198

Goodwill

5968

Other non-current assets

5069

Total assets

34622

Notes payable & short-term borrowings

1463

Accounts payable

14816

Employee compensation & benefits

1136

Taxes on earnings

340

Deferred revenue

-

Other accrued taxes

982

Accrued warranty

673

Accrued deferred revenue

1095

Accrued sales & marketing programs

2758

Other accrued liabilities

1868

Other accrued liabilities

7376

Total current liabilities

25131

U.S. dollar global notes

4647

Other long-term debt, including capital lease obligations

487

Fair value adjustment related to hedged debt

-28

Less: unamortized debt issuance cost

-17

Less: current portion of long-term debt

565

Long-term debt

4524

Pension, post-retirement, & post-employment liabilities

1645

Deferred tax liability

100

Tax liability

2063

Deferred revenue

1005

Other non-current liabilities

793

Other non-current liabilities

5606

Common stock

16

Additional paid in capital

663

Retained earnings (deficit)

-473

Net unrealized gain (loss) on available-for-sale securities

5

Net unrealized gain (loss) on cash flow hedges

291

Unrealized components of defined benefit plans

-1141

Accumulated other comprehensive income (loss)

-845

Total Hewlett-Packard Co. stockholders' equity (deficit)

-639

Total stockholders' equity (deficit)

-639

How do I find the debt to equity ratio if the total liabilities is not available
Answer #1

Total liabilities = total assets-total stock holder equity

= 34622-(-639) = 35261

debt/ equity = 35261/(-639) = -55.181